OpenStudy (anonymous):

Will Fan and Medal! Thank you♥ 1) What is the values of the function y = -2x - 4 for x = 0,1,2 and 3? a. 0, -6, -8, -10 b. -4, -6, -8, -10 ** c. -4, -2, 0, 2 d. 0, 6, 8, 10

OpenStudy (anonymous):

@flower2468 can you help me again? please :)

OpenStudy (flower2468):

sure just give me a sec:)

OpenStudy (flower2468):

0*2-4=-4, 1*2-4=-2, 2*2-4=0, 3*2-4=2 so the answer is -4, -2, 0, 2

OpenStudy (flower2468):

so C

OpenStudy (anonymous):

@flower2468 thank you so much♥

OpenStudy (anonymous):

@flower2468 can you help me with another one?